OpenStudy (anonymous):

Quick one Estimate the value of 60.22 + 21.17 + 3 x 33.2 by using the compensation strategy.

OpenStudy (campbell_st):

60 + 20 + 100 = 180

OpenStudy (anonymous):

It's not?!

OpenStudy (anonymous):

The answer at the back of my book is 181. Also, if you don't mind, can you explain how you get the answer?

OpenStudy (anonymous):

It is 60 + 21 + 100 = 181 is the answer...

OpenStudy (anonymous):

Can you further explain it?

OpenStudy (anonymous):

I'm confused because you're like doing it using reformulation strategy.

OpenStudy (anonymous):

In Approximation in case of decimal numbers,for example if I have 50.4, then you can take it as 50 but if it is 50.5 then you will take it as 51... For 1 2 3 4 digits after decimal you will take that number only but if you have 5 6 7 8 9 after decimal then you can add 1 to that number as I did...

OpenStudy (anonymous):

But what about ''compensation strategy'' ?

OpenStudy (anonymous):

I don't know the compensation strategy...

OpenStudy (anonymous):

I can handle the question using reformulation strategy.. That's probably what you've mentioned. But when it comes to compensation strategy, I can do nothing but sitting here, using reformulation strategy again. That why I asked here to see if I can seek help. By the way, thanks for your help.

OpenStudy (anonymous):

Can you tell me somewhat about what is compensation strategy??

OpenStudy (anonymous):

The line from the book is ''Get a better estimate by making adjustment'' And the example is: Estimate 2316x5 Step 1: Replace 2316 by 2000, then 2000x5 = 10000 Step 2: In order to obtain a better estimate, replace 316 by 300. So, we have 300 x 5 =1500. Step 3: Combine the results in Steps 1 and 2. The estimate value = 10000 + 1500 = 11500.

OpenStudy (campbell_st):

estimation is a real guessing game... I just work on the most obvious... and round to the nearest 10 or possibly 5